# PseudoNeg-MAE: Self-Supervised Point Cloud Learning using Conditional Pseudo-Negative Embeddings

### v2

@@ 3 Proposed Approach - PseudoNeg-MAE
A potential shortcut for COPE is to produce identity weights for all transformations, $\Theta_{g}=I$, making the features invariant to transformations. This trivial solution undermines our goal of learning transformation-sensitive representations. To avoid this, we introduce the process of conditional generation of pseudo negatives using COPE. With pseudo negatives, we regularize COPE and prevent it from collapsing into the identity. This approach ensures that different transformations produce different projections in the embedding, maintaining the sensitivity to the transformations. Specifically, we generate pseudo negatives by applying COPE to a set of randomly sampled transformations $\{g_{r}\}$ from $\mathcal{G}$, where $\{g_{r}\}$ are different from the input transformation $g$. For each $g_{r}$, the corresponding weight $\Theta{g_{r}}$ with the COPE is computed and generated the pseudo negative embedding $\mathbf{\tilde{{z}}}_{i}^{g_{r}}$:
@@
We incorporate these pseudo negatives into our novel loss function $\mathcal{L}_{\text{cope}}$ to regularize COPE.
@@
$\mathcal{L}_{\text{cope}}=\frac{1}{N}\sum_{i=1}^{N}\log\left[\sum_{r=1}^{M}e^{-\left\|\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+e^{-\left\|\frac{\Theta_{g}\mathbf{z}_{i}}{{\|\Theta_{g}\mathbf{z}_{i}\|}}-\mathbf{z}_{i}^{+}\right\|_{2}^{2}/\tau}\right],$ (3)
@@
where $M$ is the number of pseudo negatives. This loss penalizes the model for mapping different transformations to similar embeddings, thus preventing the trivial solution. Consider the case of perfect alignment where $\mathbf{z}_{i}^{+}=\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}$. Minimizing $\mathcal{L}_{\text{cope}}$ is equivalent to optimizing $\log\left[\sum_{r=1}^{M}e^{-\left\|\frac{\Theta_{g}\mathbf{z}_{i}}{{\|\Theta_{g}\mathbf{z}_{i}\|}}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+1\right]$, which effectively maximizes the pairwise distances between $\{\Theta_{g_{r}}\mathbf{z}_{i}\}$. Achieving this requires diverse $\{\Theta_{g_{r}}\}$ from the COPE, thereby penalizing trivial invariant solutions. We define the final loss $\mathcal{L}_{\text{PseudoNegMAE}}$ as follows:
$\mathcal{L}_{\text{PseudoNegMAE}}=\mathcal{L}_{\text{align}^{*}}+\beta\mathcal{L}_{\text{cope}}+(1-\beta)\mathcal{L}_{\text{unif}}.$ (4)
@@
Specifically, $\mathcal{L}_{\text{unif}}$ encourages discriminative features and $\mathcal{L}_{\text{cope}}$ preserves the sensitivity to transformations within the feature space. This combination enables the model to learn discriminative and sensitive representations to input transformations. The loss $L_{\text{align}^{*}}$ encourages the positive $\mathbf{z}_{i}$ to be pulled towards the anchor. $\mathcal{L}_{\text{cope}}$ encourages the pseudo negatives $\mathbf{\tilde{{z}}}_{i}^{g_{r}}$ to be pushed away from the anchor $\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}$. For a given $\mathbf{z}_{i}$ all the pseudo-negatives $\{\Theta_{g_{r}}\mathbf{z}_{i}\}$ only depend on the output of the COPE $\Theta_{g_{r}}$ network. This avoids $\Theta_{g}$ collapsing into identity, preventing the undesired invariant solution. Unlike SIE [8] whose equivariance is mainly enforced through an alignment loss, a term that trivially vanishes when the predictor collapses to the identity, $\mathcal{L}_{COPE}$ conditioned with the Pseudo-Negatives reaches its maximum when predictor collapse to identity. The full proof appears in the Supplementary Materials.
@@
COPE Network Architecture: COPE generates linear weights conditioned on the transformation set ${\{g}\}$. In our work, we consider rotations as transformations and utilize quaternions to represent them as inputs to the COPE. The network architecture is illustrated in the right section of Fig. 1. First, we embed $g$ into a harmonic space $H(\cdot)$, using a set of high-frequency functions [30]. Subsequently, these embedded quaternions are processed through an MLP network to extract transformation-sensitive embeddings, denoted as $\mathbf{h}_{g}^{\prime}\in\mathbb{R}^{\frac{d}{k}}$. Independently estimating each element of $\Theta_{g}$ is computationally expensive. Therefore, we utilize shared learnable embedding vectors $\Psi\in\mathbb{R}^{\frac{d}{k}\times d}$, which are responsible for the entries of the rows in $\Theta_{g}$. We perform an element-wise multiplication of these vectors with our $h_{g}^{\prime}$, i.e., $\Gamma=h_{g}^{\prime}\odot\Psi$, where $\Gamma\in\mathbb{R}^{d\times\frac{d}{k}}$. A shared non-linear MLP network is then used to expand each column of $\Gamma$ to the dimension $d$. This process yields the linear weight $\Theta_{g}\in\mathbb{R}^{d\times d}$. Consequently, $\Theta_{g}$ is determined solely based on the input transformation.
@@
We generate pseudo negatives using COPE by sampling a set of random rotations $\{{g_{r}}\}_{r=1}^{M}$. We uniformly sample quaternions representing 3D rotations to ensure coverage over the rotation space. Specifically, for each pseudo negative, we generate a random vector $\mathbf{v}_{r}=(w_{r},x_{r},y_{r},z_{r})$ by sampling from a standard normal distribution $\mathbf{v}_{r}\sim\mathcal{N}(0,I_{4})$. We then normalize $\mathbf{v}_{r}$ to obtain a unit quaternion $g_{r}$ and ensure the scalar part is nonnegative: $g_{r}=\frac{\mathbf{v}_{r}}{|\mathbf{v}_{r}|}\cdot\text{sign}(w_{r}).$ This approach ensures that the set of quaternions ${g_{r}}$ is uniformly distributed over the space of 3D rotations. Random rotations are fed to the COPE for generating pseudo negatives as in Eq. 2.
@@
PseudoNeg-MAE Transformer Backbone: Our contrastive framework is built on the Masked AutoEncoder (MAE), combining the strengths of both predictive (MAE) and discriminative (contrastive) model training paradigms. We employ a shared Siamese encoder with a point cloud transformer backbone $f$, initialized with weights from an MAE in which there are two branches with shared weights. PseudoNeg-MAE operates on pairs of point clouds: a randomly sampled point cloud $\mathbf{x}_{i}$ and its transformed point cloud $\mathbf{x}_{i}^{+}$ with the applied transformation $g$. Following the standard transformer architecture on the point cloud, we first extract a sequence of non-overlapping point patches, which are then converted to a set of tokens.
@@
Given an input point cloud $\mathbf{x}$, we first used Furthest Point Sampling (FPS) to sample a set of $n$ center points. Then, K Nearest-Neighbour (KNN) is utilized to select $k$ neighboring points around each center point to construct $n$ point patches. Following this, patch tokens are generated using a mini-PointNet [3] and fed to the transformer encoder. However, in PseudoNeg-MAE, not all the patch tokens are processed by the encoder. A random masking strategy is applied independently to $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$, masking $m$ number of the point patches. Only visible tokens are fed to an encoder ($\mathcal{E}$). This independent masking strategy encourages the learning of representations that are robust to noise and occlusions. Then, the encoded embeddings of visible tokens from the two inputs are sent to the decoders with shared weights. An extra learnable [CLS] token is appended to each encoder output. We implement a lightweight decoder $\mathcal{D}$. The objective of the decoder is to aggregate the encoder outputs into representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$.
Relative Pose Estimation Using COPE: To tackle the challenge of relative pose estimation between source and target point clouds, we introduce a novel algorithm that utilizes the COPE network to optimize the transformation parameters directly. The algorithm focuses on estimating the quaternions to minimize the difference between the transformed source and target embeddings, thus achieving precise alignment.
The process, detailed in Algorithm 1, starts by exploring a set of randomly generated quaternions $\{g_{r}\}$, representing potential rotations between the source and target. Each quaternion is evaluated in a bi-directional manner—applying the transformation to both the source and the target embeddings via $g_{r}$ and $g_{r}^{-1}$. Here $g_{r}^{-1}$ is the inverse of $g_{r}$. The quaternions are refined through an iterative process, where each is adjusted according to the gradient update weighted by $\epsilon=0.01$. After updating the quaternions, it is essential to ensure they remain valid unit quaternions, which is achieved by the normalization step at line 1. The quaternion associated with the minimum loss is designated as $g_{est}$ serving as the final solution for the alignment. This process ensures that the iterative refinement systematically approaches the global minimum of the loss landscape. In summary, this algorithm extends the capabilities of COPE to solve relative pose estimation by optimizing quaternions to align global embeddings of point clouds.
Data: Source and target $\mathbf{z}_{src},\mathbf{z}_{tgt}$
1 Initialize $\mathcal{L}_{\min}=\infty$;
2 for $j\!=\!1\ldots N_{\text{out}}$ do
3       Initialise $g_{r}=(w_{r},x_{r},y_{r},z_{r})$;
4       for $k\!=\!1\ldots N_{\text{in}}$ do
5             $\mathbf{z}^{g_{r}}\!=\!\frac{\text{COPE}(g_{r})\mathbf{z}_{src}}{\|\text{COPE}(g_{r})\mathbf{z}_{src}\|}\;$;
6             $\mathbf{z}^{g_{r}^{-1}}\!=\!\frac{\text{COPE}(g_{r}^{-1})\mathbf{z}_{tgt}}{\|\text{COPE}(g_{r}^{-1})\mathbf{z}_{tgt}\|}\;$;
7             $\mathcal{L}_{g_{r}}\!=\!\|\mathbf{z}_{tgt}-\mathbf{z}^{g_{r}}\|_{2}^{2}+\|\mathbf{z}_{src}-\mathbf{z}^{g_{r}^{-1}}\|_{2}^{2}$;
8             $g_{r}\leftarrow g_{r}-\epsilon\nabla_{g_{r}}\mathcal{L}_{g_{r}}$;
9             $g_{r}\leftarrow\displaystyle\frac{g_{r}}{\|g_{r}\|}\,\text{sign}(w_{r})$;
10
11       end for
12      if $\mathcal{L}_{g_{r}}<\mathcal{L}_{\min}$ then
13             $\mathcal{L}_{\min}\leftarrow\mathcal{L}_{g_{r}}$,$g_{\text{est}}\leftarrow g_{r}$;
14
15       end if
16
17 end for
return $g_{\text{est}}$
@@ 13             $\mathcal{L}_{\min}\leftarrow\mathcal{L}_{g_{r}}$,$g_{\text{est}}\leftarrow g_{r}$;
Algorithm 1 Transformation Estimation with COPE.
@@
Table 1: Shape classification results on ModelNet40 dataset and ScanObjNN’s OBJ-BG subset under z/z and S0(3)/SO(3) evaluation scenario.
Method ModelNet40 ScanObjNN
@@
Self-Supervised Learning

### v3

@@ III-A Sensitivity to Transformation through Pseudo-Negatives
A potential shortcut for COPE is to produce identity weights for all transformations, $\Theta_{g}=I$, making the features invariant to transformations. This trivial solution undermines our goal of learning transformation-sensitive representations. To avoid this, we introduce the process of conditional generation of pseudo-negatives using COPE. With pseudo-negatives, we regularize COPE and prevent it from collapsing into the identity. This approach ensures that different transformations produce different projections in the embedding, maintaining the sensitivity to the transformations. Specifically, we generate pseudo-negatives by applying COPE to a set of randomly sampled transformations $\{g_{r}\}$ from $\mathcal{G}$, where $\{g_{r}\}$ are different from the input transformation $g$. For each $g_{r}$, the corresponding weight $\Theta{g_{r}}$ with the COPE is computed and generated the pseudo-negative embedding $\mathbf{\tilde{{z}}}_{i}^{g_{r}}$:
@@
We incorporate these pseudo-negatives into our novel loss function $\mathcal{L}_{\text{cope}}$ to regularize COPE.
@@
$\begin{split}\mathcal{L}_{\text{cope}}&=\frac{1}{N}\sum_{i=1}^{N}\log\Biggl[\sum_{r=1}^{M}e^{-\left\|\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+\\ &\qquad\qquad\quad e^{-\left\|\frac{\Theta_{g}\mathbf{z}_{i}}{{\|\Theta_{g}\mathbf{z}_{i}\|}}-\mathbf{z}_{i}^{+}\right\|_{2}^{2}/\tau}\Biggr],\end{split}$ (3)
@@
where $M$ is the number of pseudo-negatives. This loss penalizes the model for mapping different transformations to similar embeddings, thus preventing the trivial solution. Consider the case of perfect alignment where $\mathbf{z}_{i}^{+}=\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}$. Minimizing $\mathcal{L}_{\text{cope}}$ is equivalent to optimizing $\log\left[\sum_{r=1}^{M}e^{-\left\|\frac{\Theta_{g}\mathbf{z}_{i}}{{\|\Theta_{g}\mathbf{z}_{i}\|}}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+1\right]$, which effectively maximizes the pairwise distances between $\{\Theta_{g_{r}}\mathbf{z}_{i}\}$. Achieving this requires diverse $\{\Theta_{g_{r}}\}$ from the COPE, thereby penalizing trivial invariant solutions. We define the final loss $\mathcal{L}_{\text{Point-PNG}}$ as follows:
$\mathcal{L}_{\text{Point-PNG}}=\mathcal{L}_{\text{align}^{*}}+\beta\mathcal{L}_{\text{cope}}+(1-\beta)\mathcal{L}_{\text{unif}}.$ (4)
Figure 2: Conceptual visualization of the loss $\scriptstyle\mathcal{L}_{\text{Point-PNG}{}}$. Using COPE, we generate the anchor $\scriptstyle\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}$ and a set of pseudo-negatives $\scriptstyle\frac{\Theta_{g_{r}}\mathbf{z}_{i}}{\|\Theta_{g_{r}}\mathbf{z}_{i}\|}$ for the corresponding positive $\scriptstyle\mathbf{z}_{i}^{+}$. As per our main goal, due to these pseudo-negatives, the model remains sensitive to different transformations rather than becoming completely invariant to them.
@@
Specifically, $\mathcal{L}_{\text{unif}}$ encourages discriminative features and $\mathcal{L}_{\text{cope}}$ preserves the sensitivity to transformations within the feature space. This combination enables the model to learn discriminative and sensitive representations to input transformations. Fig. 2 illustrates the objective of our loss in terms of the anchor, the positive, and the pseudo-negatives. The loss $L_{\text{align}^{*}}$ encourages the positive $\mathbf{z}_{i}$ to be pulled towards the anchor. $\mathcal{L}_{\text{cope}}$ encourages the pseudo-negatives $\mathbf{\tilde{{z}}}_{i}^{g_{r}}$ to be pushed away from the anchor $\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}$. For a given $\mathbf{z}_{i}$ all the pseudo-negatives $\{\frac{\Theta_{g_{r}}\mathbf{z}_{i}}{\|\Theta_{g_{r}}\mathbf{z}_{i}\|}\}$ only depend on the output of the COPE $\Theta_{g_{r}}$ network. This avoids $\Theta_{g}$ collapsing into identity, preventing the undesired invariant solution. Unlike SIE [10] whose equivariance is mainly enforced through an alignment loss, a term that trivially vanishes when the predictor collapses to the identity, $\mathcal{L}_{COPE}$ conditioned with the Pseudo-Negatives reaches its maximum when predictor collapse to identity. The full proof appears in the Appendix.
III-B COPE Network Architecture
@@
COPE generates linear weights conditioned on the transformation set $\{g$}\}. In our work, we consider rotations as transformations and utilize quaternions to represent them as inputs to the COPE. The network architecture is illustrated in the right section of Fig. 1. First, we embed $g$ into a harmonic space $H(\cdot)$, using a set of high-frequency functions [42]. Subsequently, these embedded quaternions are processed through an MLP network to extract transformation-sensitive embeddings, denoted as $\mathbf{h}_{g}^{\prime}\in\mathbb{R}^{\frac{d}{k}}$. Independently estimating each element of $\Theta_{g}$ is computationally expensive. Therefore, we utilize shared learnable embedding vectors $\Psi\in\mathbb{R}^{\frac{d}{k}\times d}$, which are responsible for the entries of the rows in $\Theta_{g}$. We perform an element-wise multiplication of these vectors with our $h_{g}^{\prime}$, i.e., $\Gamma=h_{g}^{\prime}\odot\Psi$, where $\Gamma\in\mathbb{R}^{d\times\frac{d}{k}}$. A shared non-linear MLP network is then used to expand each column of $\Gamma$ to the dimension $d$. This process yields the linear weight $\Theta_{g}\in\mathbb{R}^{d\times d}$. Consequently, $\Theta_{g}$ is determined solely based on the input transformation. This design fundamentally differs from the hypernetwork proposed in SIE [10], which independently predicts all $d^{2}$ elements of the transformation matrix. In contrast, COPE uses the shared learnable embedding vectors to generate the columns of $\Theta_{g}$, which significantly reduces the number of trainable parameters. Furthermore, the proposed conditional pseudo-negatives generation explicitly penalizes invariant collapse, allowing us to avoid the architectural constraint of the bias-free hypernetwork required by SIE [10].
@@
We generate pseudo-negatives using COPE by sampling a set of random rotations $\{{g_{r}}\}_{r=1}^{M}$. We uniformly sample quaternions representing 3D rotations to ensure coverage over the rotation space. Specifically, for each pseudo-negative, we generate a random vector $\mathbf{v}_{r}=(w_{r},x_{r},y_{r},z_{r})$ by sampling from a standard normal distribution $\mathbf{v}_{r}\sim\mathcal{N}(0,I_{4})$. We then normalize $\mathbf{v}_{r}$ to obtain a unit quaternion $g_{r}$ and ensure the scalar part is nonnegative: $g_{r}=\frac{\mathbf{v}_{r}}{|\mathbf{v}_{r}|}\cdot\text{sign}(w_{r}).$ This approach ensures that the set of quaternions $g_{r}$} is uniformly distributed over the space of 3D rotations. Random rotations are fed to the COPE for generating pseudo-negatives as in Eq. 2.
III-C Transformer Backbone
@@
We employ a shared Siamese encoder with a point cloud transformer backbone $f$, initialized with weights from an MAE in which there are two branches with shared weights. Point-PNG operates on pairs of point clouds: a randomly sampled point cloud $\mathbf{x}_{i}$ and its transformed point cloud $\mathbf{x}_{i}^{+}$ with the applied transformation $g$. Following the standard transformer architecture on the point cloud, we first extract a sequence of non-overlapping point patches, which are then converted to a set of tokens.
@@
Given an input point cloud $\mathbf{x}$, we first used Furthest Point Sampling (FPS) to sample a set of $n$ center points. Then, K Nearest-Neighbour (KNN) is utilized to select $k$ neighboring points around each center point to construct $n$ point patches. Following this, patch tokens are generated using a mini-PointNet [3] and fed to the transformer encoder. However, in Point-PNG, not all the patch tokens are processed by the encoder. A random masking strategy is applied independently to $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$, masking $m$ number of the point patches. Only visible tokens are fed to an encoder ($\mathcal{E}$). This independent masking strategy encourages the learning of representations that are robust to noise and occlusions. Then, the encoded embeddings of visible tokens from the two inputs are sent to the decoders with shared weights. An extra learnable [CLS] token is appended to each encoder output. We implement a lightweight decoder $\mathcal{D}$. The objective of the decoder is to aggregate the encoder outputs into representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$.
Data: Learning rate $\epsilon=0.01$
Data: Source and target embeddings $\mathbf{z}_{src},\mathbf{z}_{tgt}$
Initialize $\mathcal{L}_{min}=\infty$
for j in $N_{out}$ do
   Initialize $g_{r}=(w_{r},x_{r},y_{r},z_{r})$
 for k in $N_{in}$ do
    $\mathbf{z}^{g_{r}}=\mathrm{\text{COPE}{}}(g_{r})\mathbf{z}_{src}$
    $\mathbf{z}^{g_{r}^{-1}}=\mathrm{\text{COPE}{}}(g_{r}^{-1})\mathbf{z}_{tgt}$
    $\mathcal{L}_{g_{r}}=\left\|\mathbf{z}_{tgt}-\mathbf{z}^{g_{r}}\right\|_{2}^{2}+\left\|\mathbf{z}_{src}-\mathbf{z}^{g_{r}^{-1}}\right\|_{2}^{2}$
    $g_{r}^{*}=g_{r}-\epsilon\nabla_{g_{r}}L_{g_{r}}$
    $g_{r}=\frac{g_{r}}{\|g_{r}\|}\cdot\text{sign}(w_{r}^{*})$
 if $\mathcal{L}_{g_{r}}<\mathcal{L}_{min}$ then
      Update $\mathcal{L}_{min}=\mathcal{L}_{g_{r}}$, set $g_{est}=g_{r}$
return $g_{\mathrm{est}}$
@@
Algorithm 1 Optimization of Transformation Estimation with Minimum Loss Tracking.
III-D Relative Pose Estimation Using COPE
To tackle the challenge of relative pose estimation between source and target point clouds, we introduce a novel algorithm that utilizes the COPE network to optimize the transformation parameters directly. The algorithm focuses on estimating the quaternions to minimize the difference between the transformed source and target embeddings, thus achieving precise alignment.
The process, detailed in Algorithm 1, starts by exploring a set of randomly generated quaternions $\{g_{r}\}$, representing potential rotations between the source and target. Each quaternion is evaluated in a bi-directional manner, applying the transformation to both the source and the target embeddings via $g_{r}$ and $g_{r}^{-1}$. Here $g_{r}^{-1}$ is the inverse of $g_{r}$. The quaternions are refined through an iterative process, where each is adjusted according to the gradient update weighted by $\epsilon=0.01$. After updating the quaternions, it is essential to ensure they remain valid unit quaternions, which is achieved by the normalization step at line 1. The quaternion associated with the minimum loss is designated as $g_{est}$ serving as the final solution for the alignment. This process ensures that the iterative refinement systematically approaches the global minimum of the loss landscape. In summary, this algorithm extends the capabilities of COPE to solve relative pose estimation by optimizing quaternions to align global embeddings of point clouds. Although Algorithm 1 uses an iterative approach, the computational cost remains minimal because the optimization operates directly on pre-computed embeddings. It uses the lightweight COPE network to predict the displacements, avoiding repetitive backbone inference. Furthermore, the random initializations are independent and fully parallelizable on GPUs. This design ensures computational feasibility while avoiding the large rotational errors often occurring at high initial rotations.
@@ III-D Relative Pose Estimation Using COPE
TABLE I: Shape classification results on ModelNet40 dataset and ScanObjNN’s OBJ-BG subset under z/z and SO(3)/SO(3) evaluation scenario.
Method ModelNet40 ScanObjNN–OBJ-BG
@@
Self-Supervised
PointBERT [15] 91.5 89.6 83.8